DOUBLE RESONANCE EXCITATION OF THE RUBIDIUM DIMER : THE 2 ^{1}Π_{g} STATE

ANASTASIA DROZDOVA, Department of Chemistry, Moscow State University, 119991 Moscow, Russia ; ABDUL-RAHMAN ALLOUCHE, GHASSAN WANNOUS, PATRICK CROZET and \underline{AMANDA J. ROSS}, Institut Lumière Matière, Université Lyon 1 & CNRS UMR5306, Université de Lyon, France.

We have performed a series of optical-optical double resonance experiments with one or two cw Ti:sapphire lasers, to excite the 2 ^{1}Π_{g} state of Rb_{2}, recording infrared fluorescence from 2 ^{1}Π_{g} on a Fourier transform spectrometer. Fluorescence from the lower vibrational levels of 2 ^{1}Π_{g} (T_{e} = 22069.56 cm^{−1}) is dominated by transitions to the B ^{1}Π_{u} state studied by Amiot and Vergès^{a}. Vibrational and rotational relaxation from laser-pumped levels v’ < 15 now give a rather complete description around the potential minimum of the 2 ^{1}Π_{g} state, completing the observations for 6 ≤ v ≤ 50 reported by Han et al^{b} last year. Fluorescence from v’ > 35, occurs also to the 0^{+} components of the A ^{1}Σ_{u}^{+} ∼b ^{3}Π_{u} complex. Fitting all available 2 ^{1}Π_{g} →B ^{1}Π_{u} data for ^{85}Rb_{2} and ^{85}Rb^{87}Rb (several thousand transitions) has also given an improved description of the bottom of the B ^{1}Π_{u} state potential well. The 2 ^{1}Π_{g} state correlates at long-range with Rb 5s + Rb 4d ^{2}D_{3/2} atoms^{c}, giving a dissociation energy of 1279.6 cm^{−1}. Most new data lie below v = 45, 250 cm^{−1} below this dissociation threshold.

aAmiot and Vergès, Chem. P hys. Lett. \underline{294} 91-98 (1997)

bX. Han et al, Chem. P hys. Lett. \underline{538} 1-4 (2011)

cA.-R.Allouche, M. Aubert-Frécon, J. Chem. P hys \underline{136} 37-41 (2012)